OpenStudy (anonymous):

Tell me how much to go up and over but heres the prob: f (x)= -|x+5| -3

OpenStudy (anonymous):

\[f \left( x \right)=-\left| x+5 \right|-3\] when x=-5,f(x)=-3 \[\left| x+5 \right|=x+5,if x \ge-5\] \[\left| x+5 \right|=-\left( x+5 \right) ,if x <-5\] draw the graph